OpenStudy (anonymous):

Find the vertex of the parabola by completing the square. x^2-6x+8=y

OpenStudy (kenljw):

x^2-6x+8=y y=(x-2)(x-4) This form doesn't show parabola y=(x-3)^2 -9+8=(x-3)^2-1 this form does

OpenStudy (anonymous):

the choices A. (–3,– 6) B. (3,–1) C. (–3,1) D. (–1, 3)

OpenStudy (anonymous):

(3,-1) is the vertex